OpenStudy (anonymous):

solve. log2(log3(y))=3

OpenStudy (australopithecus):

log2(log3(y)) = 3 2^log2(log3(y)) = 2^3 log3(y) = 2^3 3^log3(y) = 3^2^3 y = 3^2^3

OpenStudy (australopithecus):

y = 3^( (2)(2)(2) ) y = 3^8

OpenStudy (australopithecus):

Key rule: \[b^{\log_b(x)} = x\] example \[3^{\log_3(x)} = y\] \[55^{\log_{55}(x)} = y\]

OpenStudy (australopithecus):

hope you follow

OpenStudy (australopithecus):

oops those last two examples should equal x not y

OpenStudy (australopithecus):

sorry about the mistake

OpenStudy (anonymous):

its ok and it may b a dumb question but where did youu get the 2 from?

OpenStudy (australopithecus):

your question is I assumed \[\log_2(\log_3(y)) = 3\]

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

so to eliminate \[\log_2\] I just made put both sides of the equation exponents to 2 then I repeated this step again to eliminate \[\log_3\]

OpenStudy (anonymous):

oh ok

OpenStudy (australopithecus):

REMEMBER THE RULE \[b^{\log_b(x)} = x\]